Examine what Big-O notation is and explain its role in evaluating efficiencies of algorithms. 


1
Expert's answer
2021-10-23T13:52:54-0400

Definition:

"f(x)=O(g(x))" as "x\to \infin"

if the absolute value of  f(x) is at most a positive constant multiple of  g(x) for all sufficiently large values of x. That is, f(x)=O(g(x)) if there exists a positive real number M and a real number x0 such that

"|f(x)|\le Mg(x)" for all "x\ge x_0"


Big O notation is a mathematical notation that describes the limiting behaviour of a function when the argument tends towards a particular value or infinity. 

In computer science, big O notation is used to classify algorithms according to how their run time or space requirements grow as the input size grows. Big O notation characterizes functions according to their growth rates: different functions with the same growth rate may be represented using the same O notation.

A description of a function in terms of big O notation usually only provides an upper bound on the growth rate of the function.